OpenStudy (anonymous):

What is 3 2\5 x 2 1\3 ? Express the answer in simplest form.

OpenStudy (anonymous):

44.8

OpenStudy (anonymous):

@pooja195 am i right

pooja195 (pooja195):

It would help the user understand if you explained your answer ^_^

OpenStudy (anonymous):

oh well honestly i did it how it is shown and put that in my calculator

OpenStudy (anonymous):

i kinda confused tho

OpenStudy (anonymous):

@mathaisgarza27

pooja195 (pooja195):

Start by making the 2 fractions imporper any ideas on how to do that?

OpenStudy (anonymous):

u in k 12

OpenStudy (tkhunny):

OP, you do not want to pursue this one with Garza. It is ENTIRELY unreasonable to multiply two numbers that are around 3 and manage to get something around 44.

OpenStudy (anonymous):

:\

OpenStudy (ladymew):

Ok so first you should change the mixed number fractions into improper fractions. In order to do that take the whole number, let's start with 3 2/5, and multiply he whole number by the denominator 3x5=15 the. Add the numerator 15+2=17 now you have 17/5 2 1/3 ----> 2x3=6, 6+2=8 so 8/3 would be your next improper fraction now let's go back and input them into the equation. 17/5 x 8/3 = 17x8 / 5x3 = 136/15 =9.06 Or 17/5 = 3.4 and 8/3=2.6 3.4x2.6=9.06 There are many ways to do this but I always go with tops (numerators) divided by bottoms (denominators)

OpenStudy (ladymew):

Oops forgot to add that 9.06= 9 3/5